OpenStudy (anonymous):

Find the equation of the line through the point (2, 5) that cuts off the least area from the first quadrant. Give your answer using the form below. y − A = B(x − C)

OpenStudy (amistre64):



OpenStudy (amistre64):

y = m(x-a) + b A = (y intercept)(x intercept)/2

OpenStudy (anonymous):

y-5=(-5/2)(x-2)

OpenStudy (amistre64):

are we just blurting out answers now?

OpenStudy (anonymous):

It is fundamental knowledge that the slope of the line is -Yo/Xo

OpenStudy (amistre64):

yes, but how do we determine that -5/2 is that slope?

OpenStudy (anonymous):

I don't think students who are taking Calculus 1 need the formal proof. This is fundamental knowledge.

OpenStudy (amistre64):

the calculus is used to find the slope ... the value is in the practice of working the problem, not in the "well the textbook says this and that" ....

OpenStudy (amistre64):

.. not that the poster is around to gain any value from it tho lol

OpenStudy (anonymous):

If I had to show students how addition worked we'd just be wasting our time. There is a fine line where fundamental knowledge doesn't need formal proofs.

OpenStudy (amistre64):

if its fundamental, and the student already knows it ... then why are they asking the question?

OpenStudy (anonymous):

I have a question for you, what's 4x3?

OpenStudy (amistre64):

do you want to know? or just have an answer given to you? do you want to understand the process? or do you just want to know if i know that you know that i can tell you that the answer is 12

OpenStudy (anonymous):

By the way, your profile picture gave me the inspiration for mine. So, thanks <3

OpenStudy (amistre64):

i recall :)

OpenStudy (anonymous):

Do you get the joke behind mine hehehehehehe

OpenStudy (amistre64):

maybe, but at my age i make it a rule to only laugh at my own jokes ... since no ones else will :)

OpenStudy (anonymous):

I have a philosophical question for you. How come whenever someone asks me which way the north pole is and I point into the ground they laugh at me?

OpenStudy (amistre64):

humor is absorbed from the ether in such a fashion that they arent laughing with you, but at you

OpenStudy (anonymous):

Since you like formal proofs, wanna prove to me that the shortest distance between two points is a straight line?

OpenStudy (amistre64):

only if we can define the path between the points properly

OpenStudy (amistre64):

the shortest distance from the point i am at to the point I am at is not a straight line ....

OpenStudy (anonymous):

Prove that the shortest distance from my seat in this room to the instructor in front of me is a straight line

OpenStudy (amistre64):

how fast can you make a spitball?

OpenStudy (anonymous):

That's not a straight line that's a parabola

OpenStudy (amistre64):

straight enough for the parthenon ... all their straight lines are curved

OpenStudy (amistre64):

the shortest displacement between 2 distinct points is a straight line; the shortest distance is the minimal path that gets you there

OpenStudy (anonymous):

Very random, but do you know the name of that movie that's played during Christmas that has that kid that beats up that other kid then his mom gets mad at him?

OpenStudy (amistre64):

that wasnt a movie, its a repressed memory :)

OpenStudy (anonymous):

What's the title

OpenStudy (amistre64):

the grinch maybe? im not much on movies

OpenStudy (amistre64):

its a wonderful life?

OpenStudy (anonymous):

I don't know, that's the only scene I remember lmao xD